Solve the system by substitution.
2x + 2y = 14
y = –3x + 9
Hint: You must find both the x and the y values for full credit.

Answer by checkley75(3666)

y=-3x+9 now substitute (-3x+9) for y in the other equation & solve for x.
2x+2(-3x+9)=14
2x-6x+18=14
-4x=14-18
-4x=-4
x=-4/-4
x=1 answer.
y=-3*1+9
y=-3+9
y=6 answer.
proof
6=-3*1+9
6=-3+9
6=6